Two horizontal wires separated by a distance of 0.
Answer #1

  two long straight parallel wires separated by a distance d carrying currents I1 and I2 .

$\displaystyle{\frac{F_2}{l}}$=$\displaystyle{\frac{\mu_0I_1I_2}{2\pi d}}$

\begin{figure} \begin{center} \leavevmode \epsfxsize=2.8 in \epsfbox{/export/home/fyde/randy/figs/fig19-9.eps}\end{center}\end{figure}

the force per unit length of the wire is F/l = (4 pi x10^-7 x i1 x i2 )/( pix r)

= 2x10^-7 x15x3 / 0.5

=1.8 x10^-5 N/m.